OpenStudy (anonymous):

A bag contains 7 orange marbles, 20 green marbles, and 11 blue marbles. If two marbles are chosen randomly from the bag one at a time and put back, what is the probability that the first marble is green and the second marble is blue?

hero (hero):

There are 38 total marbles and each marble is put back after being chosen randomly, therefore, the probability the first marble is green is\[\frac{20}{38}\]The probability the second marble is blue is: \[\frac{11}{38}\] Each event is independent. What do we do with independent probabilities to find the total probability of both events?

OpenStudy (anonymous):

add them?

hero (hero):

Not add.

OpenStudy (anonymous):

multiply?

hero (hero):

Yes

OpenStudy (anonymous):

does the denominator stay the same?

hero (hero):

Simplify before multiplying

hero (hero):

When you multiply fractions you multiply the numerators, then you multiply the denominators.

hero (hero):

Only when adding do they stay the same. Reduce the fractions before multiplying

OpenStudy (anonymous):

10/19 x 11/38 ?

hero (hero):

You can reduce the fraction even more than that

hero (hero):

\[\frac{10}{19} \times \frac{19}{38} = \frac{10}{38} \times \frac{11}{19}\]

OpenStudy (anonymous):

55/361 ?